# 1]FAIR at Meta

### v1

@@ Results
The results are presented in Table 5.6. We test four values of $p$: $0.1,0.2,0.3$ and $0.4$. Our results show that the proposed training strategy makes both LLMs more effective and efficient. We first inspect the results of the Mistral-7B models. For the slow mode inference, fine-tuning the model with trace dropping and randomized training improves upon the baseline model that is directly finetuned on Aug-MATH dataset. The absolute Greedy@1 metric improved by 1.7% when $p=0.1$ (which amounts to 10% relative performance improvement), and 0.9% with $p=0.2$ and $0.3$, and 0.1% when $p=0.4$. Our models are also outperforming the baseline model for the Pass@20 metric when $p=0.1,0.2$ and $0.3$, where the absolute correct rate increases to $61.9\%$. Under both evaluation schemes, the average length of the reasoning trace goes down as $p$ goes up. Similarly, for inference in the fast mode, our models also achieve higher correct rate. A similar trend of performance improvement also holds for the Llama-3-8B models. Finally, for reader’s reference, we include the results of both the Mistral-7B and the Llama-3-8B models finetuned on the original MATH dataset. The performance of those models clearly lags behind [^2] The trace generated by models trained on MATH is much shorter, because the answers in MATH do not contain as many intermediate steps as those in Aug-MATH..
@@ 6 Conclusion
We present a simple and easy-to-implement framework for training Transformers to solve reasoning and planning tasks. We carefully probe the structure of the reasoning traces and design corresponding dropping strategies that imitate the shortcuts in human thinking process. By randomly applying the dropping strategies to training examples, the resulting model, Dualformer, can be controlled to execute in either fast or slow reasoning mode, or in auto mode where it decides the mode to engage. Dualformer achieves enhanced performance for the maze navigation tasks and the Sokoban game, while reducing the number of reasoning steps required. Remarkably, our approach is not limited to training task-specific models from scratch. We apply techniques in the same spirit to finetune LLMs to answer math questions and obtain improved performance. Last, the proposed framework also reduces computation consumption as the input sequences are shortened after trace dropping. Future work could investigate whether our approach helps models scale, and explore methodologies such as curriculum learning and hierarchical planning to adapt Dualformer for more complex tasks.

### v2

@@ Results
The results are presented in Table 5.8. We test four values of $p$: $0.1,0.2,0.3$ and $0.4$. Our results show that the proposed training strategy makes both LLMs more effective and efficient. We first inspect the results of the Mistral-7B models. For the slow mode inference, fine-tuning the model with trace dropping and randomized training improves upon the baseline model that is directly finetuned on Aug-MATH dataset. The absolute Greedy@1 metric improved by 1.7% when $p=0.1$ (which amounts to 10% relative performance improvement), and 0.9% with $p=0.2$ and $0.3$, and 0.1% when $p=0.4$.
Our models are also outperforming the baseline model for the Pass@20 metric when $p=0.1,0.2$ and $0.3$, where the absolute correct rate increases to $61.9\%$. Under both evaluation schemes, the average length of the reasoning trace goes down as $p$ goes up.
Similarly, for inference in the fast mode, our models also achieve higher correct rate. A similar trend of performance improvement also holds for the Llama-3-8B models. Finally, for reader’s reference, we include the results of both the Mistral-7B and the Llama-3-8B models finetuned on the original MATH dataset. The performance of those models clearly lags behind [^2] The trace generated by models trained on MATH is much shorter, because the answers in MATH do not contain as many intermediate steps as those in Aug-MATH..
@@ 6 Conclusion
We present a simple and easy-to-implement framework for training Transformers to solve reasoning and planning tasks. We carefully probe the structure of the reasoning traces and design corresponding dropping strategies that imitate the shortcuts in human thinking process. By randomly applying the dropping strategies to training examples, the resulting model, Dualformer, can be controlled to execute in either fast or slow reasoning mode, or in auto mode where it decides the mode to engage. Dualformer achieves enhanced performance for the maze navigation tasks and the Sokoban game, while reducing the number of reasoning steps required. Remarkably, our approach is not limited to training task-specific models from scratch. We apply techniques in the same spirit to finetune LLMs to answer math questions and obtain improved performance. Last, the proposed framework also reduces computation consumption as the input sequences are shortened after trace dropping.
Future work could investigate whether our approach helps models scale, and explore methodologies such as curriculum learning and hierarchical planning to adapt Dualformer for more complex tasks.